Grade 12th passAlgebra

Find the whole quantity if
70% of it is 14 minutes.

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let us assume the whole quantity be x,
Then,
(70/100) × (x) = 14
X = 14 × (100/70)
X = 14 × (10/7)
X = 20 minutes

Thanks